The Square of a number exceeds the number by 110. Find the number
Found 2 solutions by mananth, issacodegard:
Answer by mananth(16946)

let the number be x
square = x^2
x^2=x+110
x^2-x-110=0
x^2-11x+10x-110=0
x(x-11)+10(x-11)=0
(x+10)(x-11)=0
x= -10 OR 11

Answer by issacodegard(60)

Hi, let x be the number in question.
We know x^2=x+110
So, x^2-x-110=0
By using the quadratic formula, we get
x=11 or x=-10